# Reinterpreting Landauer conductance, solving the quantum measurement problem, grand unification

### v1

@@ 2.2 Demonstration of $\rho_{lpd}$ from $G_{\alpha\beta}$s
and $\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$ is a functional derivative. The parameter $t_{\beta}$ controles coupling of states of the sample with the states at the STM tip for which $\nu_{\beta}$ is the relevant density of states (DOS). For a commercially available STM tip $t_{\beta}$ and $\nu_{\beta}$ are just parameters, either specified by the manufacturer or to be ingeniously approximated by the user. Thus, $s_{\alpha\gamma}^{\prime}$ is the scattering matrix element for scattering from $\gamma$ to $\alpha$ when the STM tip is drawing a current given by the RHS of Eq. 55. When the STM tip is removed by $t_{\beta}\rightarrow 0$ then this scattering amplitude will be $s^{\prime}_{\alpha\gamma}=s_{\alpha\gamma}$. Coherent currents are given by $\frac{e_{0}}{h}|s^{\prime}_{\alpha\gamma}|^{2}$ and $\frac{e_{0}}{h}|s_{\alpha\gamma}|^{2}$ in the two situations.
@@
In earlier works, Eq. 55 was studied in the semi-classical regime wherein $\rho_{lpd}$ is positive definite, which limits its value and applicability in the quantum regime. The quantum regime remained unclear specially because in the quantum regime $\rho_{lpd}$, as caclculated from Eq. 56, can become negative. We will show that Eq. 55 remains valid even when $\rho_{lpd}$ becomes negative in a quantum regime. For this it is extremely important to apply wave-mechanics or von Neumann quantum mechanics for the region between the two reservoirs, that is the mesoscopic system. The key idea needed to show this is to realize the fact that in the quantum regime of 1D, 2D and 3D, $\rho_{lpd}$ cannot really be negative although sometimes it misleads one to think so. Real negativity of $\rho_{lpd}$ is a speciality of the mesoscopic regime when we encounter Fano resonances [2]. This is the primary reason why earlier works could not understand the quantum regime and to be demonstrated below.
@@
Intuitively speaking, $\rho_{lpd}$ is positive and the LHS of Eq. 55 is negative accounting for loss of coherent electrons to the earthed lead $\beta$. Therefore, any change of sign in $\rho_{lpd}$ will counter-intuitively result in a gain of coherent current due to the intervention of the earthed STM tip on a quantum state. That can be regarded as a supporting evidence for time time travel (like a quantum state getting younger with respect to coordinate time) because any intervention on a quantum state at an intermediate point, by an STM tip is supposed to only decohere the state, with measurement outcomes being random. We will clarify as to why and when such counterintuitive effects can be seen in real systems. Since LHS of Eq. 55 can be defined in quantum mechanics and can also be measured in the laboratory and so if the equality in Eq. 55 can be established then one can affirm that RHS is an objective reality that can be found in nature, although we cannot make sense of it within the axiomatic framework of quantum mechanics. The object $\rho_{lpd}$ cannot be defined within the axiomatic framework of quantum mechanics because it is refering to only those electrons that are coming from $\gamma$ and going to $\alpha$. The wave-function of an electron at a point r can only be sensitive to either $\gamma$ or $\alpha$. But the object can be defined for a mesoscopic system by appealing to the idea of a physical clock.
If we start from a closed system govorned by Schrodinger equation, we cannot get an object that is $\gamma$ and $\alpha$ dependent because in that case the idea of a physical clock do not work as they inherently need asymptotic free states [2]. In a mesoscopic set up these asymptotic states can not be in a linear superposition. The classical reservoirs destroy the linear superposition of states in the leads while linear superposition exists inside the sample (shaded region) [2]. Events in the reservoir are all classical events and quantum event can only be defined in between two classical events in the reservoirs. This will be discussed again with respect to the 1D picture given in Fig. 3. Once we know $\rho_{lpd}$ we can use it to find other relevant quantities like injectance that we find in axiomatic framework of quantum mechanics and so we are not loosing anything but only gaining some extra objects by this assertion. At any point we can go to the closed system by making the leads $\gamma$ and $\alpha$ very weakly connected to the system and recover all that we get by solving the Schrodinger equation for the closed system. Besides it may not be always necessary to pinch off the leads physically as von Neumann quantum mechanics give us a method to infer about the closed system from the scattering states as has been worked out in chapter 4 of Ref [2].
@@
In this paragraph we discuss that $\rho_{lpd}$ as a local object solves the problem of collapse of wave-function by making it redundant, with measurements defined deterministically in the classical world of the reservoirs, while the time evolution of observables in the region between the reservoirs is determined by quantum mechanical scattering. At a point r in the shaded region, there will be electrons in a linear superposition of states. One may say that all these electrons came from lead $\gamma$ (assume for the time being that $\rho_{lpd}$ is positive definite) but some may be going to lead $\alpha$ (we call them dancing cats), some may be going to lead $\beta$ (we call them sleeping cats) and some may be getting reflected back to lead $\gamma$ (we call them beaten cats). At the point r, we have a linear superposition of dancing, sleeping and beaten cats. But we can now say something about only the dancing cats. They spend a time $\frac{h}{2\pi}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$, at the point r. At zero temperature there are $|s^{\prime}_{\alpha\gamma}|^{2}$ of these dancing cats at that point and we can average over only the dancing cats to define a local partial density of states given by $\rho_{lpd}(E,\alpha,\textbf{r},\gamma)=-\frac{1}{2\pi}\;|s^{\prime}_{\alpha\gamma}|^{2}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$. That means we can selectively average over the dancing cats at the point r and determine physical effects due to the dancing cats alone. In formal quantum mechanics we cannot do this selective averaging for the dancing cats alone that are in a linear superposition with the sleeping and beaten cats. Note that $|s^{\prime}_{\alpha\gamma}|^{2}$ and $\theta_{s^{\prime}_{\alpha\gamma}}$ are measured at the junction of the sink reservoir which is a classical system. Probability is well defined for classical systems and measured $|s^{\prime}_{\alpha\gamma}|^{2}$ is a probability wherein measurements do not care how the probability is defined theoretically. Similarly, according to the notion of physical clock, $\theta_{s^{\prime}_{\alpha\gamma}}$ is defined as angular displacement of a precessing dipole. A perfectly logical connection can be made between quantum mechanical scattering phase shift $\theta_{s^{\prime}_{\alpha\gamma}}$ and well defined angular displacement of a precessing dipole [2]. Thus the measurables are not random outcomes induced by collapse of wavefunction by the intevening STM tip at r, but can be calculated precisely in the Landauer-Buttiker formalism wherein the punchline is that a quantum event can only be defined between two classical events in the reservoirs that perfectly help define a local time and a local state at the point r. This will be demonstrated with the help of Eq. 55 wherein the LHS has the measurables and the RHS has the local object. We know that von Neumann quantum mechanics has one mathematical framework for scattering states and bound states. For any scattering phenomenon the LPDOS can be defined and bound states can be seen as a special case of the scattering states. So we believe that this solves the measurement problem in general and we have to see how the Landauer-Buttiker formalism can be extended to other practical cases. At various places in this work, we have sketched out this roadmap as to what is a general mesoscopic system and how any other system can be seen as a modification of it. That is why we have discussed von Neumann formalism, Diracs approach, Heisenberg’s approach, S. Datta’s approach, reservoirs being in the domain of relativity, etc.
@@
Thus we intend to justify Eq. 55 for the quantum regime, specially the situation when $\rho_{lpd}$ becomes negative. Ref. [8] gives a derivation of Eq. 55, we will give a physical topological arguement here which will make it clear why and when we can generally expect the equality in Eq. 55 inspite of $\rho_{lpd}$ being negative. That will be followed by a numerical demonstration. Note that there are two DOS that feature in Eq. 55, that are $\nu_{\beta}$ and an r dependent DOS that is $\rho_{lpd}$. There is a factor $\frac{1}{2\pi}$ for each of them that cancels the factor $4\pi^{2}$. This is the $\frac{1}{2\pi}$ factor that appears in Eq. 56. There can be a very simple interpretation for this factor as follows. LHS of Eq. 56 is a DOS that can accommodate electrons that are countable but $2\pi\rho_{lpd}$ is uncountable, the RHS of Eq. 56 being in terms of $\theta_{s_{\alpha\gamma}}(E)$ corresponding to a continuous rotation. This also make it essential that if only one half of a Riemann surface is involved then rotaion angle can only be maximum $\pi$ instead of $2\pi$. So winding number in Eq. 56 as a conserved quantity in the two cases will differ by a factor of 2. On the other hand, scattering probabilities $|s^{\prime}_{\alpha\gamma}|^{2}$ and $|s_{\alpha\gamma}|^{2}$ are just numbers, are well defined in quantum mechanics, and when multiplied by a factor $\frac{e_{0}}{h}$ they give the measured coherent current from $\gamma$ to $\alpha$.
@@
Let us consider the system shown in Fig. 3. First of all it is a system for which exact quantum mechanical calculations can be made and it is a system that exhibit pronounced Fano resonances. The system in Fig. 2 will generally show such resonances was shown in earlier works. In this sense the system in Fig. 2 at Fano resonances with the modification that the upper lead is weakly coupled to earth via a strong potential can be represented by the 1D system in Fig. 3. Besides, since the logic, to be presented below, is based on topology of the Argand diagram (AD) at a resonance and the complex plane, there are only two kinds of behavior, one being typically like that of Breit-Wigner resonance that involve the full Riemann surface, and the other being typically like that of Fano resonance that involve one half of the Riemann surface. Also since it is a question of whether LPDOS is present at all in a quantum system and more importantly whether its negativity is allowed in quantum mechanics, then demonstrations in 1D quantum mechanics should suffice. Let us drop the factor $|t|^{2}\nu_{\beta}$ from Eq. 55 because such factors will be only needed if we are using a commercially available STM tip, where this factor characterizes the STM tip. This require us to drop a $2\pi$ (or $\pi$) factor in Eq. 55 for reasons mentioned in the previous paragraph. Which means for a first principle calculation for the system in Fig. 3 we expect (if the equality in Eq. 55 is correct) that
@@
In Fig. 4 we plot the Argand diagram (AD) for $t_{31}$ for the system shown in Fig. 3 by varying the tunneling potential $U_{1}$ which connects the three prong potential with lead 2 that is earthed. As $U_{1}$ is varied from $e_{0}U_{1}l=-10$ to $e_{0}U_{1}l=-1000$ we plot $Im[t_{31}]$ versus $Re[t_{31}]$, which is known as AD. The AD diagram make smooth sub-loops within one Riemann surface and each sub-loop is due to a Fano resonance. Any physical quantity that depend on $t_{31}$ will go through a cycle over one particular sub-loop. For one closed sub-loop generated by monotonously varying a parameter, say $U_{1}(\textbf{r})$, the local potential at r.
$\oint_{c}\Delta\theta_{t_{31}}=\oint_{c}\frac{\delta\theta_{t_{31}}}{\delta U_{1}({\bf{r}})}\;\Delta U_{1}({\bf{r}})=0\;\;\mbox{where}\;\;\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$
Because the AD was generated by varying $U_{1}(\textbf{r})$, the integration measure is $\Delta U_{1}(\textbf{r})$ and the integration for a closed loop will be zero. Or
@@
Therefore from Eqs. 60 and 61 we get the equality
@@
purely as a consequence of the topology of a complex plane $(Im[t_{31}],Re[t_{31}])$. In this complex plane $(Im[t_{31}],Re[t_{31}])$, the point (0,0) is a singular point where the phase $\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$ is undefined and is well known for classical waves [9]. If the Argand diagram trajectory would have enclosed the singularity, then the RHS of Eq. 60 would have been $2\pi$ instead of zero but the RHS of Eq. 61 would be still zero, in which case Eq. 62 will not be valid. We get Eq. 62 from Eqs. 60 and 61 because due to Fano resonances, one gets loops in the AD within one Riemann surface. Whatever be the terms in the Hamiltonian or the equation of motion, it is the topology of the complex plane that determines the outcome of the integration in Eqs. 60 and 61. For the kind of Argand diagram in Fig. 4, that are smoothly closed within one Riemann surface, the equality in Eq. 62 is for the integrals. We will address the value of the integrands latter but the integrands will go through a positive-negative cycle over a loop simply because of the fact that the integration over a loop will be zero and there is no surprize if $\rho_{lpd}$, which is the integrand in Eq. 60, becoming negative. However, the question remains if it is physical and if it can be found in nature. That can be verified by also considering the integrand in the LHS of Eq. 62.
Therefore, it is obvious from Fig. 4 and Eq. 62, that $\Delta|t_{31}|^{2}$ as well as $\rho_{lpd}$ will go through a positive-negative cycle over a closed loop as a consequence of the topology of the relevant complex plane, only if the AD closes within one side of the Riemann surface, as in the case of Fig 4, the AD is restricted to the first and the fourth quadrant. Given the fact that cycles in $\Delta|t_{31}|$ is physical, the cycles of $\rho_{lpd}$ is also physical as they are just different expressions of the same loop made by the AD of $t_{31}$. If we want to compare the integrands in Eq. 62, then we need an extra factor of $2\pi$ as in Eq. 58 and 59 as derived in [8]. Thus
@@
where primed quantities and unprimed quantities are calculated for an infinitesimal difference of $U_{1}$. If we want to compare integrands then we can only write an approximate equality because we are comparing two different objects over one sub-loop. There will always be in the least a phase difference between the two quantities. But if $\rho_{lpd}$ determines the measured quantity $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ then the two sides of Eq. 64 should agree in amplitude and frequency. The agreement in frequency will definitely be there because that is determined by the same AD.
@@
In Fig. 5 we plot LHS and RHS of Eq. 64 where the primed values and unprimed values are again for small differences in $U_{1}$ as the incident wave vector $k$ is varied and that is the horizontal axis. Initially the two curves are a little different but as $k$ increases, effects of dispersion is minimized, and magnitudes of the two curves are similar with a phase difference between them. The oscillations of both the curves are due to the smooth cyclic nature of the AD in Fig 4. In other words, the underlying principle behind the oscillation of both curves is the same. The two curves oscillate as a result of quantum interference. But the symmetric oscillation between positive and negative values and the similarity between the two curves can only be noted if there are well defined Fano resonances and signify the fact that the RHS of Eqs. 58 and 59 are physical. For Breit-Wigner resonances when the AD starts going around the singular point there will be deviations. Since Buttiker et al did not consider Fano resonances, the negativity of $\rho_{lpd}$ is unphysical and they did not understand Eqs. 58 and 59 in the quantum regime.
@@
Figure 5: Fig. 5. In this figure we are plotting the LHS and RHS of Eq. 64 to show that they both can oscillate between positive and negative values. The primed and unprimed values are for small differences in $U_{1}$ at a $k$ value that we vary continuously. The sign change and magnitude of both the curves originate from the smooth cyclic AD of Fig. 4.

### v2

@@ 2.2 Demonstration of $\rho_{lpd}$ from $G_{\alpha\beta}$s
and $\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$ is a functional derivative. The parameter $t_{\beta}$ controls coupling of states of the sample with the states at the STM tip for which $\nu_{\beta}$ is the relevant density of states (DOS). For a commercially available STM tip $t_{\beta}$ and $\nu_{\beta}$ are just parameters, either specified by the manufacturer or to be ingeniously approximated by the user. Thus, $s_{\alpha\gamma}^{\prime}$ is the scattering matrix element for scattering from $\gamma$ to $\alpha$ when the STM tip is drawing a current given by the RHS of Eq. 55. When the STM tip is removed by $t_{\beta}\rightarrow 0$ then this scattering amplitude will be $s^{\prime}_{\alpha\gamma}=s_{\alpha\gamma}$. Coherent currents are given by $\frac{e_{0}}{h}|s^{\prime}_{\alpha\gamma}|^{2}$ and $\frac{e_{0}}{h}|s_{\alpha\gamma}|^{2}$ in the two situations.
@@
In earlier works, Eq. 55 was studied in the semi-classical regime wherein $\rho_{lpd}$ is positive definite, which limits its value and applicability in the quantum regime. The quantum regime remained unclear specially because in the quantum regime $\rho_{lpd}$, as caclculated from Eq. 56, can become negative. We will show that Eq. 55 remains valid even when $\rho_{lpd}$ becomes negative in a quantum regime. For this it is extremely important to apply wave-mechanics or von Neumann quantum mechanics for the region between the two reservoirs, that is the mesoscopic system. The key idea needed to show this is to realize the fact that in the quantum regime of 1D, 2D and 3D, $\rho_{lpd}$ cannot really be negative although sometimes it misleads one to think so. Real negativity of $\rho_{lpd}$ is a specialty of the mesoscopic regime when we encounter Fano resonances [2]. This is the primary reason why earlier works could not understand the quantum regime and to be demonstrated below.
@@
Intuitively speaking, $\rho_{lpd}$ is positive and the LHS of Eq. 55 is negative accounting for loss of coherent electrons to the earthed lead $\beta$. Therefore, any change of sign in $\rho_{lpd}$ will counter-intuitively result in a gain of coherent current due to the intervention of the earthed STM tip on a quantum state. That can be regarded as a supporting evidence for time travel (like a quantum state getting younger with respect to coordinate time) because any intervention on a quantum state at an intermediate point, by an STM tip is supposed to only decohere the state. We will clarify as to why and when such counterintuitive effects can be seen in real systems. Since LHS of Eq. 55 can be defined in quantum mechanics and can also be measured in the laboratory and so if the equality in Eq. 55 can be established then one can affirm that RHS is an objective reality that can be found in nature, although we cannot make sense of it within the axiomatic framework of quantum mechanics.
@@
In this paragraph we discuss that $\rho_{lpd}$ as a local object solves the problem of collapse of wave-function by making it redundant, with measurements defined deterministically in the classical world of the reservoirs, while the time evolution of observables in the region between the reservoirs is determined by quantum mechanical scattering. At a point r in the shaded region, there will be electrons in a linear superposition of states. One may say that all these electrons came from lead $\gamma$ (assume for the time being that $\rho_{lpd}$ is positive definite) but some may be going to lead $\alpha$ (we call them dancing cats), some may be going to lead $\beta$ (we call them sleeping cats) and some may be getting reflected back to lead $\gamma$ (we call them beaten cats). At the point r, we have a linear superposition of dancing, sleeping and beaten cats. In formal quantum mechanics, we cannot look at the cats at the point r and say which one is dancing (meaning going to lead $\alpha$) etc. But we can now say something about only the dancing cats. They spend a time $\frac{h}{2\pi}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$, at the point r. At zero temperature there are $|s^{\prime}_{\alpha\gamma}|^{2}$ of these dancing cats at that point and we can average over only the dancing cats to define a local partial density of states given by $\rho_{lpd}(E,\alpha,\textbf{r},\gamma)=-\frac{1}{2\pi}\;|s^{\prime}_{\alpha\gamma}|^{2}\;\frac{\delta\theta_{s^{\prime}_{\alpha\gamma}}}{\delta U({\bf{r}})}$. That means we can selectively average over the dancing cats at the point r and determine physical effects due to the dancing cats alone. Like for example the LHS of Eq. 55 is physical and is determined by the dancing cats, only thing that remains to be verified is that the equality in Eq. 55 holds. In formal quantum mechanics we cannot do this selective averaging for the dancing cats alone that are in a linear superposition with the sleeping and beaten cats. Note that $|s^{\prime}_{\alpha\gamma}|^{2}$ and $\theta_{s^{\prime}_{\alpha\gamma}}$ are measured at the junction of the sink reservoir which is a classical system. Probability is well defined for classical systems and measured $|s^{\prime}_{\alpha\gamma}|^{2}$ is a probability wherein measurements do not care how the probability is defined theoretically. Similarly, according to the notion of physical clock, $\theta_{s^{\prime}_{\alpha\gamma}}$ is defined as angular displacement of a precessing dipole. A perfectly logical connection can be made between quantum mechanical scattering phase shift $\theta_{s^{\prime}_{\alpha\gamma}}$ and well defined angular displacement of a precessing dipole [2]. Thus the measurables are not random outcomes induced by collapse of wavefunction by the intervening STM tip at r, but can be calculated precisely in the Landauer-Buttiker formalism wherein the punchline is that a quantum event can only be defined between two classical events in the reservoirs that perfectly help define a local time and a local state at the point r. We know that von Neumann quantum mechanics has one mathematical framework for scattering states and bound states. For any scattering phenomenon the LPDOS can be defined and bound states can be seen as a special case of the scattering states. So we believe that this solves the measurement problem in general and we have to see how the Landauer-Buttiker formalism can be extended to other practical cases. At various places in this work, we have sketched out this roadmap as to what is a general mesoscopic system and how any other system can be seen as a modification of it. That is why we have discussed von Neumann formalism, Diracs approach, Heisenberg’s approach, S. Datta’s approach, reservoirs being in the domain of relativity, etc.
@@
Thus we intend to justify Eq. 55 for the quantum regime, specially the situation when $\rho_{lpd}$ becomes negative. Ref. [8] gives a derivation of Eq. 55, we will give a physical topological argument here which will make it clear why and when we can generally expect the equality in Eq. 55 in spite of $\rho_{lpd}$ being negative. That will be followed by a numerical demonstration. Note that there are two DOS that feature in Eq. 55, that are $\nu_{\beta}$ and an r dependent DOS that is $\rho_{lpd}$. There is a factor $\frac{1}{2\pi}$ for each of them that cancels the factor $4\pi^{2}$. This is the $\frac{1}{2\pi}$ factor that appears in Eq. 56. There can be a very simple interpretation for this factor as follows. LHS of Eq. 56 is a DOS that can accommodate electrons that are countable but $2\pi\rho_{lpd}$ is uncountable, the RHS of Eq. 56 being in terms of $\theta_{s_{\alpha\gamma}}(E)$ corresponding to a continuous rotation. This also make it essential that if only one half of a Riemann surface is involved then rotaion angle can only be maximum $\pi$ instead of $2\pi$. So winding number in Eq. 56 as a conserved quantity in the two cases will differ by a factor of 2. On the other hand, scattering probabilities $|s^{\prime}_{\alpha\gamma}|^{2}$ and $|s_{\alpha\gamma}|^{2}$ are just numbers, are well defined in quantum mechanics, and when multiplied by a factor $\frac{e_{0}}{h}$ they give the measured coherent current from $\gamma$ to $\alpha$.
@@
Let us consider the system shown in Fig. 3. First of all it is a system for which exact quantum mechanical calculations can be made and it is a system that exhibit pronounced Fano resonances. The system in Fig. 2 will generally show such resonances was shown in earlier works [2]. In this sense the system in Fig. 2 at Fano resonances with the modification that the upper lead is weakly coupled to earth via a strong potential can be represented by the 1D system in Fig. 3. Besides, since the logic, to be presented below, is based on topology of the Argand diagram (AD) at a resonance and the complex plane, there are only two kinds of behavior, one being typically like that of Breit-Wigner resonance that involve the full Riemann surface, and the other being typically like that of Fano resonance that involve one half of the Riemann surface. Also since it is a question of whether LPDOS is present at all in a quantum system and more importantly whether its negativity is allowed in quantum mechanics, then demonstrations in 1D quantum mechanics should suffice. Let us drop the factor $|t_{\beta}|^{2}\nu_{\beta}$ from Eq. 55 because such factors will be only needed if we are using a commercially available STM tip, where this factor characterizes the STM tip. This require us to drop a $2\pi$ factor in Eq. 55 for reasons mentioned in the previous paragraph. Which means for a first principle calculation for the system in Fig. 3 we expect (if the equality in Eq. 55 is correct) that
@@
In Fig. 4 we plot the Argand diagram (AD) for $t_{31}$ for the system shown in Fig. 3 by varying the tunneling potential $U_{1}$ which connects the three prong potential with lead 2 that is earthed. As $U_{1}$ is varied from $e_{0}U_{1}l=-10$ to $e_{0}U_{1}l=-1000$ we plot $Im[t_{31}]$ versus $Re[t_{31}]$, which is known as AD. The AD diagram make smooth sub-loops within one Riemann surface and each sub-loop is due to a Fano resonance, because they do not go around the origin (0,0). Any physical quantity that depend on $t_{31}$ will go through a cycle over one particular sub-loop. For one closed sub-loop generated by monotonously varying a parameter, say $U_{1}(\textbf{r})$, the local potential at r.
$\oint_{c}\Delta\theta_{t_{31}}=\oint_{c}\frac{\delta\theta_{t_{31}}}{\delta U_{1}({\bf{r}})}\;\Delta U_{1}({\bf{r}})=0\;\;\mbox{where}\;\;\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$ (60)
Because the AD was generated by varying $U_{1}(\textbf{r})$, the integration measure is $\Delta U_{1}(\textbf{r})$ and the integration for a closed loop will be zero simply because the AD does not go around the origin. Or
@@
Therefore from Eqs. 61 and 62 we get the equality
@@
purely as a consequence of the topology of a complex plane $(Im[t_{31}],Re[t_{31}])$. In this complex plane $(Im[t_{31}],Re[t_{31}])$, the point (0,0) is a singular point where the phase $\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$ is undefined and is well known for classical waves [9]. If the Argand diagram trajectory would have enclosed the singularity, then the RHS of Eq. 60 being a winding number, would have been $2\pi$ instead of zero but the RHS of Eq. 62 would be still zero, in which case Eq. 63 will not be valid. We get Eq. 63 from Eqs. 61 and 62 because due to Fano resonances, one gets loops in the AD within one Riemann surface. Whatever be the terms in the Hamiltonian or the equation of motion, it is the topology of the complex plane that determines the outcome of the integration in Eqs. 61 and 62. For the kind of Argand diagram in Fig. 4, that are smoothly closed within one Riemann surface, the equality in Eq. 63 is for the integrals. We will address the value of the integrands latter but the integrands will go through a positive-negative cycle over a loop simply because of the fact that the integration over a loop will be zero and there is no surprise if $\rho_{lpd}$, which is the integrand in Eq. 61, becomes negative. However, the question remains if it is physical and if it can be found in nature. That can be verified by also considering its equality with the integrand in the LHS of Eq. 63 as demands Eq. 59. Just to remind once again Ref. [3] could not make sense of the negativity of $\rho_{lpd}$, because they did not consider Fano resonances.
Therefore, it is obvious from Fig. 4 and Eq. 63, that $\Delta|t_{31}|^{2}$ as well as $\rho_{lpd}$ will go through a positive-negative cycle over a closed loop as a consequence of the topology of the relevant complex plane, only if the AD closes within one side of the Riemann surface, as in the case of Fig 4, the AD is restricted to the first and the fourth quadrants. However, cycles are similar only means qualitative agreement, we have to show they are quantitatively equal to claim the equality in Eq 59, or to claim that $\rho_{lpd}$ is physical, or to claim that the local states determine measurement outcomes. If we want to compare the integrands in Eq. 63, then we need an extra factor of $2\pi$ as in Eq. 58 and 59 as derived in [8]. Thus
@@
where primed quantities and unprimed quantities are calculated for an infinitesimal difference of $U_{1}$. If we want to compare integrands then we can only write an approximate equality because we are comparing two different objects over one sub-loop. There will always be in the least a phase difference between the two quantities. But if $\rho_{lpd}$ determines the measured quantity $|t^{\prime}_{31}|^{2}-|t_{31}|^{2}$ then the two sides of Eq. 65 should agree in amplitude.
@@
In Fig. 5 we plot LHS (solid curve) and RHS (dashed curve) of Eq. 65 where the primed values and unprimed values are again for small differences in $U_{1}$ as the incident wave vector $k$ is varied and that is the horizontal axis. Initially the two curves are a little different but as $k$ increases, effects of dispersion is minimized, and magnitudes of the two curves are similar with a phase difference between them. The oscillations of both the curves are due to the smooth cyclic nature of the AD in Fig 4. In other words, the underlying principle behind the oscillation of both curves is the same. The two curves oscillate as a result of quantum interference. But the symmetric oscillation between positive and negative values and the similarity between the two curves can only be noted if there are well defined Fano resonances and signify the fact that the RHS of Eqs. 58 and 59 are physical.
Note that for Fig. 5 we have set $l_{2}=l_{3}=0$ which implies that we have only Fano resonances [2]. In case of Fig. 6 we introduce a finite value for them, that is $l_{2}=l_{3}=l_{5}=1$ which means there is an interplay of Fano and Breit-Wigner type resonances. However, the equality of Eq. 65 is still qualitatively and quantitatively valid except that the dashed curve is slightly pushed up with respect to the solid curve, the reason for which will be clear from the next case. Although there are only two possibilities, that are, the AD remain restricted to one Riemann surface without enclosing the origin (as in the case of Fig. 4) or the AD goes around the origin and goes to higher Riemann surface (as in the case of Fig. 7), there interplay can lead to diversity. For example, in Fig. 7 we show the AD for a set of parameters. In this case if we follow the solid curve, the AD is initially restricted to one Riemann surface, starts from point A and gets very close to the origin at the point B but does not enclose the origin, and then goes through the points CDEFG enclosing the origin. At the point F it comes back to the point C making CDEF a closed loop encircling the origin but not coming back smoothly as in the case of Fig. 4 and instead forming a cusp at C or F. For the choice of parameters in Fig. 7, if we try to test the equality in Eq. 65 then we show in Fig. 8 that we still have very good quantitative agreement but qualitatively the LHS of Eq. 65 (solid curve) is very different from the RHS of Eq. 65 (dashed curve). Such qualitative disagreement completely overshadows the quantitative agreement between the solid and dashed curves in Fig. 8 and that can be confusing. However, if we look at the topology of the complex plane $\bigg(Im[t_{31}],Re[t_{31}]\bigg)$ then it is again possible to read off the solid curve by looking at the dashed curve. For that one has to notice that as we move along the AD by varying the energy (or wave-vector) going through the points CDEF then $\theta_{t_{31}}=Arctan\frac{Im[t_{31}]}{Re[t_{31}]}$ changes monotonously to make a complete $2\pi$ rotation. We can see from Eq. 7 that varying energy and varying electrostatic potential in some region has similar effects, specially because winding number is a topological invariant. In other words as energy is varied or $U_{1}$ is varied, $\theta_{t_{31}}$, changes monotonously and so the integration in Eq. 61 will not be zero and equality of the integrands in Eq. 63 cannot be expected. Eq. 62 will still be valid and so while the solid curve in Fig. 8 oscillates only on the positive side, the dashed curve oscillates on both sides. Some of the peaks in the solid curve, are very close to the peaks in the dashed curve but other peaks are reversed in sign. Never the less both the curves are expressing the same Argand diagram and on careful examination of how the AD behave in the $(Im[t_{31}],Re[t_{31}])$ plane with respect to the point (0,0) we can infer the dashed curve is physical meaning LPDOS is physical. Experimentalists can in fact directly observe the AD [10] that can help us further to deduce one curve from the other. That $\theta_{t_{31}}$, changes monotonously with $U_{1}$ is shown in Fig. 9. Since Buttiker et al did not consider Fano resonances, away from the Fano resonances the negativity of $\rho_{lpd}$ is unphysical and they did not understand Eq. 58 and 59 for the quantum regime.
@@
Figure 5: Fig. 5. In this figure we are plotting the LHS and RHS of Eq. 65 to show that they both can oscillate identically between positive and negative values. The primed and unprimed values are for small differences in $U_{1}$ at a $k$ value that we vary continuously. The sign change and magnitude of both the curves originate from the smooth cyclic AD of Fig. 4.
Figure 6: Fig. 6. In this figure we are plotting the LHS and RHS of Eq. 65 for a different choice of parameters for the system in Fig. 4. The primed and unprimed values are for small differences in $U_{1}$ at a $k$ value that we vary continuously.
Figure 7: Fig. 7. In this figure we are plotting the AD as wave-vector is varied in a certain range for two values of $U_{1}$ with other parameters fixed and shown in the inset.
Figure 8: Fig. 8. In this figure we are plotting the LHS of Eq. 65 (solid curve) and the RHS of the same equation in dashed curve. The primed and unprimed values are for small differences in $U_{1}$ at a $k$ value that we vary continuously.
Figure 9: Fig. 9. In this figure we are showing that $\theta_{t_{31}}$, changes monotonously with $U_{1}$ between the solid and dashed curves in Fig. 7. Thus, since $\theta_{t_{31}}$ changes monotonously with $U_{1}$ as well as with wave-vector, then $(\theta_{t^{\prime}_{31}}-\theta_{t_{31}})$ can only vary on one side of the horizontal axis. Real part of $t_{31}$ or imaginary part of $t_{31}$ can change sign as wave-vector is varied but $\theta_{t_{31}}$ changes monotonously with $U_{1}$ in the wave-vector range considered in Fig. 8.
Figure 10: Fig. 10 AD as $U_{1}$ is varied.